Please solve the value of X for:
log3 4x (l0g 4x to the base 3) = log6 x ( log x to hte base 6)

Answer by stanbon(75887)

Please solve the value of X for:
log3 4x (l0g 4x to the base 3) = log6 x ( log x to hte base 6)
----
log3(4x) = log6(x)
---
Convert:
log(4x)/log(3) = log(x)/log(6)
---
log(6)*log(4x) = log(3)*log(x)
---
0.7782*log(4x) = 0.4771*log(x)
0.7782*log(4) + 0.7782*log(x) = 0.4771*log(x)
0.3011*log(x) = -0.4685
log(x) = --1.5560
---
x = 10^-1.5560
x = 0.0278
================
Cheers,
Stan H.